I need help with this question on my homework. The answer I have gotten is 7/45. Please help me solve this word problem: If a committee of five is to be chosen from six males and eight females, what is the probability that the committee will be all male or all female?
Answer by stanbon(75887)

If a committee of five is to be chosen from six males and eight females, what is the probability that the committee will be all male or all female?
----
P(5 males) = 6C5/14C5 = 6/2002
-----
P(5 females) = 8C5/14C5 = 56/2002
=====================================
Cheers,
Stan H.
=========================